OpenStudy (anonymous):

The following identitit invovle the reciprocal, quotient, and Pythagorean relationships. Prove each one. cos^2(x)cos^2(y) + sin^2(x)sin^2(y) +sin^2(x)cos^2(y) + sin ^2(y)cos^2

OpenStudy (primeralph):

That's not an identity. That's an expression.

OpenStudy (anonymous):

I copied word for word what is on the sheet, I belive i have to break this down into simplier terms

OpenStudy (primeralph):

It doesn't equal anything?

OpenStudy (anonymous):

Is there no way to simplify this?

OpenStudy (primeralph):

You can simplify it, but you can't prove anything.

OpenStudy (anonymous):

How would I simplify it?

OpenStudy (primeralph):

The last part is incomplete.

OpenStudy (anonymous):

that is how it is written

OpenStudy (primeralph):

sin ^2(y)cos^2 Okay then. Good luck.

OpenStudy (raden):

the last part is sin ^2(y)cos^2(x) or sin ^2(y)cos^2(y) ???

OpenStudy (anonymous):

idk it just ended as sin^2(y) cos^2 <- it is missing x or y, possible typo on teachers part?

OpenStudy (raden):

i guess the last part is sin ^2(y)cos^2(x). if so, we can simply it be : cos^2(x)cos^2(y) + sin^2(x)sin^2(y) +sin^2(x)cos^2(y) + sin ^2(y)cos^2(x) = ( cos^2(x)cos^2(y) + sin ^2(y)cos^2(x) ) + ( sin^2(x)sin^2(y) +sin^2(x)cos^2(y) ) = cos^2(x) (cos^2(y) + sin^2(y)) + sin^2(x) (sin^2(y) + cos^2(y)) = cos^2(x) * 1 + sin^2(x) * 1 = cos^2(x) + sin^2(x) = 1

OpenStudy (raden):

for the 1st step, it like a+b+c+d. it would be (a+d) + (b+c)

OpenStudy (anonymous):

you grouped them?

OpenStudy (raden):

yes, then factor out which the common's and use the identity sin^2 + cos^2 = 1

OpenStudy (anonymous):

alright I worked it out, is there any way to remember which steps to take? because after each question I have a hard time just starting :c

OpenStudy (primeralph):

Factorize, recognize identities, simplify.